OpenStudy (anonymous):

Solve for x in the proportion 3 times x over the quantity 3 times x minus 36 equals negative 1 over x. x = 3 and x = −4 x = −3 and x = 4 x = −3 x = −12

OpenStudy (dumbcow):

\[\frac{3x}{3x-36} = \frac{-1}{x}\] \[\frac{3x}{3(x-12)} = \frac{-1}{x}\] \[\frac{x}{x-12}=\frac{-1}{x}\] cross-multiply \[x^{2} =-(x-12)\] set equal to 0 \[x^{2}+x-12 = 0\] Factor \[(x+4)(x-3) = 0\]

OpenStudy (anonymous):

whick one is the answer im trying to finish this online class

OpenStudy (dumbcow):

what makes (x+4) = 0 ?? what makes (x-3) =0 ??